OpenStudy (anonymous):

how do you arrange 1/(2x^2) for integration?

OpenStudy (anonymous):

\[\frac{1}{2}\int x^{-2}dx\]would be a good first step

OpenStudy (campbell_st):

its just \[\frac{1}{2} x^{-2}\]

OpenStudy (anonymous):

thanks, got it